40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 6842     Entry time: Thu Jun 21 01:58:29 2012
Author: Jenne 
Type: Update 
Category: Green Locking 
Subject: Xgreen preparations 

[Yuta, Koji, Jenne]

Lots of small things happened tonight, in preparation for having both arms' ALS working simultaneously.

1. Xarm aligned in IR

     1.1 ETMX oplev centered

2. Xgreen coarsely aligned to Xarm

3. X beat setup on PSL table resurrected. 

     3.1 Steering optics for both X and Y green (before PBS) were touched to fix clipping Xgreen on some of the first mirrors after the light exits the chambers.

     3.2 Xgreen aligned to beat PD

     3.3 PSL green waveplate rotated so ~half of the light goes to X beat, other ~half goes to Y beat (recall we had rotated the polarization so we had max light on the Y beat PD a few weeks ago).

          3.3.1 Now we have ~80uW of PSL green going to each beat PD.

     3.4 PSL green aligned to X beat PD

          3.4.1 Replaced mount for mirror between PBS (which splits PSL green light) and BS (which combines PSL green and X green) so that I could get the alignment correct without having to use the full range of the knobs on the mount.

     3.5 Realigned (coarsely) Ygreen to Y beat PD - the mirrors just after the chambers had been touched, so Y green was no longer directly on the PD.  This will need to be done more finely when we're ready to lock the Yarm again.

     3.6 Dedicated cables for the DC of each beat PD were put in place, so we have those in addition to the DC transmission PDs which we are putting in temporarily each time we align the green to the cavities.  Some mystery unused cables that were running under the PSL table were removed.  The power for the X beat PD was rerouted so that it's much closer to the actual diode, and out of the way. 

4. Better alignment of X green to X arm.

     4.1 Put Green Transmission camera into place

     4.2 Noticed that the X green spot on the transmission camera is not nearly as steady as the Y green.  Increased the gain of the X green refl PD on the end table to see if it helped the spot be more steady, but it's still very wiggly.  We reverted the gain to what it was.  We need to fix this!!!!

    4.3 Removed camera, looked at X transmission DC (PD is temporarily in front of the beat PD), tried to increase the transmission.

     4.4 Aligning the green to the X arm has been really tough - there were a few more iterations of camera then DC PD.

     4.5 Measured X green power on the PSL table - 02 mode was ~150uW.  The 00 mode is still not very stable, which is frustrating, although we have a reasonable amount of power transmitted.   

     4.6 The X end green shutter was moved out of the beam path since the green beam was clipping while going through the shutter.  We need to put it back now that the beam is pretty much aligned.  The beam size and the aperture are roughly the same, so we should look to see if there is a different place on the table where the beam is a little smaller, where we can put the shutter.

5. Whitening filters (Pomona box-style) made for the Xarm I and Q channels - these are the same as the whitening for the Y arm.

6. 30m SMA cable made to be used for 2nd delay line.

     6.1 Steve reminded me this morning that we returned one of the fancy spools of cable that was purchased for the delay lines, since it was defective.  We didn't get it replaced because there was debate as to what is the best kind of cable to use.  We need to come to a conclusion, but for now we have a regular RG-405 cable.

7.  Jamie has started work on modifying the beatbox so that we can have 2-arm ALS.  Hopefully that will be done soon-ish, because we're otherwise pretty close to being ready.

ELOG V3.1.3-